OpenStudy (anonymous):

:) Who wants an easy medal??? Express 5^3 = x as a logarithmic equation. log(5)3 = x log(3)5 = x log(3)x = 5 log(5)x = 3 I just totally blanked.

OpenStudy (anonymous):

You pull the exponent out front in the log rule

OpenStudy (anonymous):

\[\large b^x=y\iff \log_b(y)=x\]

OpenStudy (anonymous):

yes, what @satellite73 said

OpenStudy (anonymous):

Soo.... log(3)5=x ????

OpenStudy (anonymous):

\[5^3=125\iff \log_5(125)=3\] in other words

OpenStudy (anonymous):

I'm confused...So it's log(5)x=3

OpenStudy (anonymous):

did you understand anything about the log rules?

OpenStudy (anonymous):

its been a while, i kinda forgot

OpenStudy (anonymous):

if you have 3^2 = 9 so it is log3(9) = 2

OpenStudy (anonymous):

log(b)y=x right?

OpenStudy (anonymous):

the number on the outside is the base and then the value with your exponent at the end

OpenStudy (anonymous):

5^3=x, 5 is the base so you have log5(x)=3

OpenStudy (anonymous):

hope that helps out

OpenStudy (anonymous):

yea thank you :)